### v2

# An Empirical Study of Cross-Lingual Transfer Learning in Programming Languages

Abstract
Large language models (LLMs) have become remarkably good at improving developer productivity for high-resource programming languages. These models use two kinds of data: large amounts of unlabeled code samples for pre-training and relatively smaller amounts of labeled code samples for fine-tuning or in-context learning. Unfortunately, many programming languages are low-resource, lacking labeled samples for most tasks and often even lacking unlabeled samples. Therefore, users of low-resource languages (e.g., legacy or new languages) miss out on the benefits of LLMs. Cross-lingual transfer uses data from a source language to improve model performance on a target language. It has been well-studied for natural languages, but has received little attention for programming languages. This paper reports extensive experiments on four tasks using a transformer-based LLM and 11 to 41 programming languages to explore the following questions. First, how well does cross-lingual transfer work for a given task across different language pairs. Second, given a task and target language, how should one choose a source language. Third, which characteristics of a language pair are predictive of transfer performance, and how does that depend on the given task. Our empirical study with 1,808 experiments reveals practical and scientific insights, such as Kotlin and JavaScript being the most transferable source languages and different tasks relying on substantially different features. Overall, we find that learning transfers well across several programming languages.
Figure 1: Overview. Consider an LLM pretrained on unlabeled code in multiple seen languages. Finetune on task-specific labeled samples from a source language. For RQ1, test performance on a target language. For RQ2, train a model that ranks transfer performance given features of a language pair. For RQ3, measure how important the features of language pairs are. Repeat for several languages and four tasks: Error Detection, Solution Domain Classification, Clone Detection, and Code Repair.
I Introduction
Large language models (LLMs) leverage the naturalness of software [1] to achieve state-of-the-art performance on many software engineering tasks such as error detection, clone detection, and code translation [2, 3]. However, LLMs have thus far been trained and evaluated mainly on programming languages with vast amounts of openly available code rather than low-resource ones. Programming languages used for critical and strategic applications may have much code only in private repositories and are thus ignored. COBOL, which supports many critical financial applications and may have over 775 billion lines of code overall [4], is low-resource due to lack of open availability. Similarly with Fortran, which is used in high performance and scientific computing domains. Notwithstanding their popularity [5, 6], new languages such as Rust and Swift start out low-resource, by definition. Other languages such as R may be low-resource for licensing reasons [7]. Besides being underrepresented in training of LLMs, organizations often spend significant resources on training software developers to work with unfamiliar low-resource languages. This adds to the cost of maintaining such software and may also stifle innovation [8]. There is a need for state-of-the-art AI tools to augment the capabilities of developers for these low-resource languages. Indeed, LLMs are being used in tools to enhance developer productivity and can also help with migration or modernization of projects from one programming language to another by supporting code translation [9, 10].
Recent works have shown the possibility of transfer learning: leveraging data from one programming language to compensate for the lack of data in the target programming language [11, 12, 13]. Another recent work has developed several similarity metrics to decide which high-resource programming language dataset can be used to augment fine-tuning data for some other language [14]. Prior work, however, has used at most 6 programming languages, none of which are truly low-resource. Further, the notion of similarity among programming languages is underexamined: more research is needed to make reliable claims. To overcome these shortcomings, we examine transferability from source to target language combinations for 4 tasks and for many more languages than prior work (11–41 depending on the task). The four tasks we focus on are clone detection, code repair, solution domain classification, and error detection. Fig. 1 gives a high-level overview of the methodology for our extensive empirical study. From left to right, we start with a pretrained LLM that has seen unlabeled code in several programming languages. Next, for each of four tasks, for each source language with a suitable amount of labeled training data for that task, we finetune the LLM. For each target language with a suitable amount of labeled test data for the task, we evaluate the finetuned LLM. For example, Fig. 1 illustrates this with a solution domain classification task, using C as the source language and Rust as the target language, which yields an F1 score of $0.48$. Doing this for four different tasks and all their source and target language combinations yields four different heat maps.
Our first research question (RQ1) asks how learning transfers across programming languages in general, by exploring these heat maps. Our second research question (RQ2) asks whether one can predict the ranking of source languages in the heat maps from the previous experiment just using features of language pairs, by training a performance prediction model from the language pair features and the ground-truth target labels from the heat maps. For example, in Fig. 1, the performance prediction model for the solution domain classification task might rank C behind Kotlin as a source language for transfer to Rust. Given the cost of training LLMs, the performance prediction model is useful in directing data acquisition efforts and deciding how to spend compute resources. Finally, as our third research question (RQ3), we seek software-engineering insights into the characteristics of programming languages that affect transferability. To do this, we measure the importance of language pair features in the prediction model from the previous experiment.
This paper makes the following contributions:
  • Evaluating the pairwise transferability for several programming languages (including low-resource ones) and four tasks using LLMs.
  • Developing a method to identify the best language to transfer from, for different target languages and tasks.
  • Characterizing the features of programming language pairs that are predictive of transferability for given tasks.
We are planning to make our data generation, training, and analysis code available. One goal of this paper is to offer practical guidance to engineers who build LLM-driven software engineering tools towards more informed choices for data acquisition and modeling. A second goal of this paper is to help advance software engineering as a science by using the lens of transfer learning to shed light on programming language characteristics. Below are some results highlights:
  • Learning transfers well for all tasks. More specifically, cross-lingual learning transfers better than zero-shot.
  • Transfer learning depends on the target language. Java, Go, Dart, and TypeScript are among the best target languages. C++, Python, Vim, and COBOL are among the worst.
  • Transfer learning also depends on source language used for finetuning. Kotlin and JavaScript are the best source languages, C++ is the worst.
  • We can predict which source language can perform well for a given task by considering a variety of features.
  • Different tasks rely on substantially different features.
  • Overall the most important features are the overlap of keywords and names between the source and target languages.
If we take the premise that software engineers will increasingly benefit from LLM support, then we hope this paper will help democratize such support into low-resource settings.
II Background and Related Work
Cross-lingual Transfer for Natural Languages
Lin et al. [15] use 4 tasks to study transfer among up to 60 natural languages and explore how important features of language pairs are for predicting how well learning transfers between them, finding that feature importances vary a lot across tasks. Lauscher et al. [16] use 5 tasks to study transfer among up to 15 languages and argue that while zero-shot performance can work for low-level tasks, higher-level tasks benefit from at least a few target-language shots. de Vries et al. [17] use 1 task (POS tagging) with 65 source and 105 target languages and study the effects of language families and writing systems, with Romanian emerging as a particularly good source language. Ahuja et al. [18] use 11 tasks with 1 source and varying numbers of target languages, confirming earlier findings that feature importances vary a lot across tasks. Our paper takes inspiration from these works, especially de Vries et al. [17], but differs by studying programming languages instead of natural languages, which have different tasks and different features of language pairs affecting transferability.
Cross-lingual Transfer for Programming Languages
Zhou et al. [5] call for studying transfer, motivating with 1 task (code completion) and 2 languages (from Hack to Rust). Chen et al. [14] use 2 tasks (code summarization and search) and 6 languages to study transfer to Ruby, and propose picking the source language based on language similarity. Ahmed et al. [11] use 3 tasks to study transfer among 6 languages, demonstrating that due to the nature of their tasks, signals from identifiers are highly important for transferability. Yuan et al. [12] use 1 task (automated program repair) to study transfer among 5 languages, sequentially fine-tuning on multiple languages with innovative tricks to prevent catastrophic forgetting. Pian et al. [13] use 2 tasks (code summarization and completion) to study transfer among 4 languages, using meta-learning to improve a base learner. Our paper also focuses on programming languages, but considers more tasks and vastly more languages to explore conditions for effective transfer.
Besides the above works, there is also work on learning for multiple programming languages. While such prior work has been instrumental in making studies like ours possible, it differs by not focusing on cross-lingual transfer. Our work benefits from datasets for multiple programming languages. CodeNet covers 55 languages [19]; XCodeEval covers 7 tasks with up to 17 languages [20]; and MultiPL-E supports 1 task (code generation) in 19 languages [21]. There are also models for multiple programming languages. TransCoder uses pre-training across 3 languages to help transpilers [10]. CodeT5 is pre-trained on 8 languages and 5 tasks [22]. StarCoder is pretrained on 86 languages [7], but e.g. does not include COBOL and evaluation is largely focused on Python. Unsupervised pretraining on multiple languages has become common, but transferability of supervised tasks has not yet been thoroughly studied; our paper addresses that gap.
III Experimental Setup
As illustrated in Fig. 1, given a task, our experimental approach first finetunes a model for each source language individually and then tests each finetuned model on all target languages. We applied this approach to four tasks where the number of source languages varies from 6 to 22, leading to 58 finetuned models in total. For each source language, we finetune the model using one A100 GPU for 6-30 hrs depending on the task. Each model was then evaluated on 11 to 43 target languages producing 1,808 experiments. All results are presented in Figs. 5(a)–(d) and analyzed in Sec. IV.
This section presents methodological details.
(a) Clone Detection: 21 source languages $\times$ 41 target languages. Metric: F1 Score.
Figure 2: Transfer scores heat map. The figure shows scores for every combination of source and target language. Each column corresponds to a source language, with “zsh” showing zero-shot performance, i.e., without fine-tuning on any source language. Each row corresponds to a target language. The languages whose language-name label uses red font were seen during pre-training. Framed black boxes highlight the performance of a source language (column) on itself as the target language (row). The dendrograms show results of hierarchical clustering based on similarity of the performance vectors. The row and column order is also determined by the same clustering.
(a) Code Repair: 20 source languages $\times$ 38 target languages. Metric: BLEU score.
Figure 3: Transfer scores heat map (cont.; see caption from 2(a))
(a) Solution Domain Classification: 11 source languages $\times$ 11 target languages. Metric: F1 score.
Figure 4: Transfer scores heat map (cont.; see caption from 2(a)).
(a) Error Detection: 6 source languages $\times$ 11 target languages. Metric: F1 Score.
Figure 5: Transfer scores heat map (cont.; see caption from 2(a)).
Base Datasets
The four tasks studied are derived from two multilingual code datasets: CodeNet [19] and XCodeEval [20]. CodeNet [19] consists of about 14M code samples for a total of 500M lines of code in 55 different programming languages, derived from submissions to online coding judge websites. The dataset comes with benchmarks for code classification and code similarity. XCodeEval [20] contains 25M samples in 11 programming languages from another different online coding judge website. This dataset comes with an execution-based evaluation framework and several different benchmarks: from classification (solution domain classification and error detection) to generation (program synthesis, automatic program repair, and code translation).
Tasks
To study how learning transfers between programming languages, we explore two type of tasks: classification (error detection, solution domain classification, clone detection) and generation (code repair). For the classification tasks, we follow Lewis et al. [23] by predicting labels from the vocabulary of class labels based on the final decoder hidden state. We use BLEU score [^1] We chose BLEU score because it enables experiments with the larger number of languages in CodeNet [19], and because other prominent benchmarks for the code repair task also adopt this metric [24]. for evaluating the generative task and F1 score for evaluating the classification tasks.
  • Error Detection: Given a code $C$ in language $L$, the task is binary classification of whether or not $C$ compiles (or can be loaded by an interpreter) without error.
  • Solution Domain Classification: Given a code $C$ in language $L$, the task is multi-label classification of $C$ into a set of tags corresponding to potential algorithmic techniques required to write the solution (e.g., sorting, graphs).
  • Clone Detection: Given two code samples $C_{1}$ and $C_{2}$ in language $L$, the task is binary classification of whether they are type-IV clones (semantically similar) [25]. To generate the dataset, we apply the following procedure to the CodeNet dataset. Given all combinations of solutions to all problems in language $L$, we identify positive samples (clones) as pairs of accepted solutions for the same problem and the others as negative examples. To balance the positive samples across problems, we ensure a ratio of 0.15 of positive samples across different languages.
  • Code Repair: Given a buggy code $C$ in language $L$, the task is to generate the corresponding repaired code. We generate the dataset by using the following procedure on CodeNet solutions: Each sample is modified by sequentially inserting, removing, or replacing tokens of different types for fixed ratios for different token types.
Data Sampling
Given variation in the size of datasets for different languages, we follow the sampling procedure of de Vries et al. [17] to select the number of training samples and identify potential source languages. We first finetune the model with datasets of different number of samples $N$={10K, 30K, 50K, 70K, 100K}. Depending on the number of training examples $N_{L}$ for a language $L$, we randomly upsample languages with $N<N_{L}$ or downsample languages with $N>N_{L}$. We filter the set of source languages for a given task based on the relative performance compared to a baseline model. We select sample sizes of 50K for solution domain classification, 70K for error detection, and 100K for clone detection and code repair. Fig. 8 in the supplementary material shows how the metric for each task vaires with the sample size.
Model
Our experimental framework is based on CodeT5-base (220M parameters) leveraging the Hugging Face transformers library [26]. Code-T5 is an open-source model, pre-trained on eight programming languages (Ruby, JavaScript, Go, Python, Java, PHP, C, and C#). Due to its encoder-decoder nature, the model performs well on both code generation and code understanding tasks [22] like code repair, error detection, and clone detection, among others. This, along with its relatively small size, makes it a good fit for our empirical study, which requires 58 finetuned models with numerous inference runs each on four tasks.
IV Results and Discussion
This section addresses the following research questions.
RQ1:
How well does cross-lingual transfer work for a given task across different language pairs?
RQ2:
Given a task and target language, how should we pick the source language for best performance?
RQ3:
Which characteristics of a language pair are predictive of transfer performance, and how does that depend on the given task?
IV-A Transfer analysis
We explore RQ1 via extensive experiments on every combination of source and target programming language, for each task. The four heat maps of Fig. 5 show the results. A source programming language is the language used for fine-tuning a model, with samples in the training data. A target programming language is the language used to evaluate a model, with samples in the test data. Fig. 5 shows source languages on the X-axis and target languages on the Y-axis. Monolingual scores (where the source and the target language are the same) are highlighted with bold boxes.
For RQ1 we analyze model performance trends with respect to task, source language, and target language. We provide some practical insights, like which languages are the toughest target languages and which are the best source languages. We also discuss some interesting source-and-target language pairs.
Task dependency
The heat maps in Fig. 5 show that transferability of source languages varies depending on the task. Visually this can be seen by how the color gradient varies for each task. To quantify this observation, we calculate the mean scores across all the language combinations. The mean cross-lingual scores are 0.78 for clone detection, 0.75 for code repair, 0.65 for error detection, and 0.47 for solution domain classification. The monolingual scores (in bold boxes) are higher (more red) than the cross-lingual scores, averaging 0.91 for clone detection, 0.98 for code repair, 0.78 for error detection, and 0.51 for solution domain classification. The zero-shot columns, where the model is not finetuned on any source language, have lower scores (more blue) for all heat maps. The mean zero-shot scores are 0.49 for clone detection, 0.28 for code repair, 0.49 for error detection, and 0.01 for solution domain classification. Variations in the mean monolingual, cross-lingual, and zero-shot scores across tasks confirm that model performance is task dependent. The difference in overall zero-shot and cross-lingual scores indicates that in the absence of fine-tuning data for any task in any programming language, it is better to finetune with some other language in that task than using a zero-shot setting. This cross-lingual advantage over the zero-shot setting occurs for every task.
Looking at these scores in absolute terms reveals that the choice of tasks also affects transferability. For solution domain classification, the difference in mean cross-lingual and monolingual scores is only 0.04, whereas for code repair, it is 0.23. The difference in mean cross-lingual and zero-shot scores for the same two tasks is 0.46 and 0.48. Although overall trends show learning transfers well in all tasks, cross-lingual performance in relation to monolingual performance is highly task dependent. Within each task, model performance depends on the choice of source and target programming language, which we explore next. Tab. II in the supplementary material shows the mean score by task.
Target language dependency
Target programming languages fall into two categories, high- and low-resource. In Fig. 5, high-resource target programming languages are represented on both axes. Low-resource target programming languages are represented only on the Y-axis. Models are finetuned only on high-resource programming languages and tested on both high- and low-resource programming languages.
To understand overall trends across all four tasks, we consider ten target programming languages that are common to all four tasks. Six (Kotlin, JavaScript, Java, Python, C, and C++) are high-resource in all four tasks. Four (Go, Rust, PHP, and C#) are low-resource in error detection. For each of these ten languages, for each of the four tasks, we calculate the mean cross-lingual score using all scores in the row of Fig. 5 for each language, excluding the monolingual score in the bold box and the zero-shot score in the extreme left. We exclude monolingual and zero-shot scores so as to understand transferability from different languages to the target language. Based on this mean we rank each of the 10 languages for each task, and then calculate the mean rank for each language.
We find Java, Go, Rust, JavaScript, and Kotlin are target languages that benefit most from transfer learning, while PHP, C#, C, C++, and Python benefit least, with Java, Go being the best and Python, C++ being the worst. C++ has a reputation as being one of the hardest programming languages to learn for human software developers, which may explain our finding that it is a difficult target for LLM transfer learning. Some target programming languages are good at monolingual performance, but much worse at cross-lingual performance; one example across multiple tasks is C++. In-language performance for C++ is 0.95, 0.99, and 0.46 for clone detection, code repair, and solution domain classification, whereas its mean cross-language performance is 0.79, 0.81, and 0.41 for the same tasks. Java is an example of more consistent in-language and cross-language transferability across tasks and hence can be considered relatively more transferable than C++. Java has in-language scores of 0.95, 0.99, and 0.45 for clone detection, code repair, and solution domain classification, and cross-language scores of 0.86, 0.84, and 0.42 for the same. All numbers are reported Tab. III in the supplementary material.
We do a similar analysis for low-resource target languages in clone detection and code repair, since these two tasks have a wider and more varied set of low-resource languages. Here we find Dart, TypeScript, Elixir, Lua, and Swift benefit most from cross-lingual training, with Dart being the best. Dart, TypeScript have average scores of 0.87 and 0.79, 0.82 and 0.83 for clone detection and code repair. COBOL, F#, Octave, bc, and Vim are least likely to benefit, with Vim being the worst. Vim has average scores of 0.69, 0.67 and COBOL has average scores of 0.72, 0.74 for Clone Detection and Code Repair. Tab. IV in the supplementary material has more information on mean rank and average score of low-resource target languages for clone detection and code repair.
For error detection, only a smaller set of languages can be considered low resource. We observe low variation in average scores across different sources with highest being Rust with 0.66 and lowest being C# with 0.64. For solution domain classification there are no low-resource target languages.
Our analysis shows that transferability depends on the choice of task and target language. We can see by comparing with zero-shot scores that all target programming languages benefit from cross-lingual training, however some languages benefit much more than others.
Most Transferable Source Language
The number of source languages vary for each of the four tasks. To understand the source language dependency and to identify the most transferrable source language, we start from languages all four tasks have in common (JavaScript, C, Kotlin, Java, C++, and Python). For each task, for each language, we calulate the mean cross-lingual score. As an example, consider the column for source language Python on the X-axis of Fig. 5(a). We calculate the mean over the scores except 0.94, which is the monolingual score with target language Python. Once we have the cross-lingual mean for each language, we rank the source languages based on their mean score for each task. Based on these rankings, we calculate the mean rank, for overall cross-lingual performance of source languages, across all target languages, across all tasks.
The most transferrable source languages from best to worst are Kotlin, JavaScript, Java, Python, C, and C++. It may be surprising that Kotlin is the best-performing source language, since it is rarely discussed in the context of training LLMs and is not part of the pre-training corpus for CodeT5 [22]. Prior work on transfer learning for natural languages [17] yielded a similar surprise, with Romanian as the best-performing source language. A possible explanation could be that Romanian and Kotlin both have strong roots in dominant languages of the past (Latin and Java), while sitting at a cross-roads that helped them absorb many influences. Detailed scores and rankings are in Tab. V in the supplementary material.
The above analysis considers all target languages, including high-resource target languages. In a more realistic setting, we can ignore high-resource target languages because the monolingual option is available for every high-resource language in every task. If we repeat the above analysis with only low-resource languages, we are left with a smaller set of target languages and only three tasks (all target languages for solution domain classification are high-resource). For the remaining three tasks and six common source languages, analysis shows that the most transferrable source languages are JavaScript, Java, Kotlin, C, Python, and C++, with JavaScript being the best and C++ being the worst. JavaScript and Java contribute the most data to CodeT5 pretraining; Kotlin and C++ are absent [22]. For details see Tab. VI in the supplementary material.
If we consider only clone detection and code repair, we get a much broader set of about 20 common source languages. Repeating the analysis for these two tasks for low-resource target languages only, we find once again that JavaScript is the most transferable and C++ is among the least. In this case the most transferable languages are JavaScript, Scala, and Haskell, while the least transferable are Fortran, AWK, C++, Lisp, and Pascal. More details are in Tab. VII in the supplementary material.
Language Pair Performance
The best model performance for each task is almost always in the monolingual setting. However, some cross-lingual pairs do stand out and show very good model performance, comparable to monolingual. Another reason to consider individual language pairs is that from the perspective of developers working with low-resource languages, it would be helpful to know which programming languages act as the best source language.
For clone detection in Fig. 5(a), notice that a model finetuned on Ruby does almost as well on Crystal, a language whose syntax is inspired by Ruby. Fig. 5(a-b) also shows that surprisingly, Haskell does very well on COBOL, much better than FORTRAN does on COBOL. From the same heat maps notice Lisp does very well on Scheme, which can be expected as Scheme is a dialect of Lisp. Similary Python does very well on Cython. Fig. 5(a-d) also show that C-category languages (C, C++, and C#) do well on each other. Details are in Tab. VIII in the supplementary material.
Dendrograms
Dendrograms for each task cluster programming languages with similar transfer learning performances. In the code repair task, syntaxically and semantically related languages are next to each other. For example, C is paired with C++ and is next to the pair C# and Java. JavaScript is paired with TypeScript, Python with Cython, Lua with Moon, Lisp with Scheme. On the other hand, relatedness of languages plays a smaller role in the dendrograms for other tasks. This might be because while the code repair task requires deep language understanding, other tasks depend more upon other features. For instance, Ahmed and Devanbu [11] have shown that for some tasks, identifiers are more important. Section IV-C will analyze the relationship between features of language pairs and their transfer performance.
1. Learning transfers well for all tasks. Cross-lingual learning transfers better than zero-shot. 2. Learning depends on the target language. Among low resource languages, Dart and TypeScript are the best while Vim and COBOL are among the worst. 3. Kotlin and JavaScript are among the best target languages, C++ is among the worst. 4. Some cross-lingual language pairs show particularly good results, such as Ruby and Crystal, Haskell and COBOL.
IV-B Performance Prediction
Feature Measure Description Linguistic Features Object oriented e Is the language object oriented? Type strength e Is the language strongly or weakly typed? Type checking e Is the language typed statically or dynamically? Type safety e Is the language type safe? Garbage collection e Does the language use garbage collector? Standardized e Is the language standardized by a committee? Expression of types e Are the types written explicitly? Paradigm o Paradigms supported (e.g., functional, imperative) Type compatibility o Is language nominally-typed or structurally-typed? Parameter passing o How are params. passed (e.g. by value, by name)? Syntactic Features Name o Number of names which overlap Text o Number of text data which overlap Keyword o Number of keywords which overlap Literal o Number of literals which overlap Punctuation o Number of punctuation signs which overlap Operator o Number of operators which overlap Comment o Number of comment tokens which overlap Syntax o Number of AST nodes which overlap Tokens o Number of tokens which overlap Dataset-Specific Features Difficulty $x$ Average difficulty of dataset problems Length $x$ Average number of tokens Time limit $x$ Average time limit of dataset problems Memory limit $x$ Average memory limit of dataset problems Model-Specific Features Pretrained s Source language is included during pretraining Pretrained t Target language is included during pretraining Pretrained b Both source and target languages in pretraining
TABLE I: Features of language pairs. Column “Measure” indicates how the feature is computed: (e)quivalence, (o)verlap, (s)ource, (t)arget, (b)oth source and target, (rd) for relative difference. For dataset specific features, ($x$) can be s, t, or rd. Details of the linguistic features are available in supplementary material Tab. IX
Given a task and target language, how should we pick the source language for best performance? Knowing this without having to train many models on different source languages can save significant resources. It also provides a basis to study important features that characterize a successful transfer from a language to another. To predict the performance of the models on the different tasks for different source languages, we train a ranking model based on gradient-boosted decision trees (GBDT) [27]. We show that our models outperform baselines based on regressors or programming-language history. We consider the features defined in Tab. I. These features can be grouped into four categories.
  1. 1.
    Linguistic features: general properties characterizing a language pair like whether they support the same paradigms, the same style of memory management (is there a garbage collector or not), or the same type-system characteristics.
  2. 2.
    Syntactic features: properties of the syntax of a language pair measured by overlap in the counts of certain token types, as determined using the Pygments library on the corpus of code in the given languages.
  3. 3.
    Dataset-specific features: properties of the problems associated with code samples for the language pair in the online coding judge website from which the dataset came.
  4. 4.
    Model-specific features: whether the model saw the source language, target language, or both during pretraining.
To predict the top source languages for a given task and target language $L_{t}$ in the set of target languages $T$, we train a ranker using the LightGBM [27] implementation of the LambdaRank algorithm [28]. The model takes features in Tab. I as inputs and scores source languages $L_{s}$ in terms of their relevance to the target $L_{t}$. Our ranker uses a boosting ensemble of 100 decision trees with 16 leaves each. We consider the normalized discounted cumulative gain score (NDCG@K) at $K=3$ as our evaluation metric. We evaluate the model using leave-one-out (LOO) cross-validation on the set $T$. For each target language $L_{t}$, we train a ranker to predict rankings of different sources for each language in $T$ leaving out the source ranking for $L_{t}$ as a test set. For each fold, we compute the NDCG@3 score on the test set. We compare the performance of our ranker with a regression-based ranker and a history ranker. The history ranker selects the top $K$ source languages with the shortest path to the given target language in a programming languages history graph [^2] http://rigaux.org/language-study/diagram.html (contracted to merge different versions of the same language). The regression-based ranker uses a LightGBM regressor with the same hyperparameter settings as our ranker and selects the top $K$ source languages with the highest predicted scores. Fig. 6 shows the mean and standard deviation of NDCG@3 using LOO on $T$. While the error bars are overlapping, both regression and rankers outperform the history ranker on all tasks, indicating the importance of considering a variety of features for predicting top-performing sources rather than simplified heuristics. Our model outperforms baseline rankers on all tasks with the exception of clone detection, for which it shows slightly lower performance than the regression model. This indicates that ranking is the best suited method for predicting top performing sources. Based on these results, we use our ranker for the feature analysis in RQ3.
Figure 6: NDCG@3 scores for different rankers and tasks corresponding to LOO evaluation over the set of target languages.
Using a ranking model and considering a variety of features rather than simplified heuristics is crucial for predicting top performing sources.
IV-C Feature analysis
Figure 7: Normalized SHAP values aggregated by tasks for the features defined in Tab. I. The features are sorted by mean rank. Grayed-out features were unavailable for the given task.
While Sec. IV-A explored how well transfer learning performs for different languages on different tasks, we would like to dig deeper into why they perform that way. That is, how do the features of language pairs affect transfer? Which characteristics of the source and target language best explain the variation of transfer across tasks, the dependence of transfer on source and target pairing, as well as the superior transfer supported by specific sources? These questions can be answered by measuring feature importance.
Unfortunately, most traditional techniques for measuring feature importance suffer from a common pitfall: when two features overlap, they arbitrarily assign more importance to one at the expense of the other. The state-of-the-art technique for measuring feature importance avoids this pitfall by considering all possible subsets $S\subseteq F$ of features [29]. Then, the importance of a given feature is the difference it makes to model performance in combination with all possible sets of other features. The resulting SHAP values [29] show the relative impact of each feature on the model output. In the context of ranking different source languages for cross-lingual transfer, SHAP values measure how different features contribute to transfer. Fig. 7 shows SHAP values computed based on the ranking models evaluated in Sec. IV-B. For presentation purposes, it rescales the values so the most important feature for each task has importance 1.
Task-dependent feature importance
While previous works emphasize the importance of specific features for cross-lingual transfer, the lack of comparison on different tasks limits their conclusions. The supported cross-task analysis demonstrates a task-dependent feature importance whereby different features contribute differently to cross-lingual transfer depending on the task. For example, the top features for the solution domain classification task are Difficulty, Literals, Names, and Operators. Being able to predict the solution domain for a code solution requires some knowledge of the underlying problem for which the tags are an attribute. The difficulty score is another attribute of the problem. Problems with similar difficulty scores require similar algorithms. The literals, names, and operators are other indicators of the algorithms used in a code sample. In comparison, the most significant features for the clone detection task are Token, Names, and Keywords. Detecting a clone requires different skills than classifying domains. A deeper understanding of the code semantics irrespective of the similar algorithms used is needed. The overlaps in names, keywords, and more generally tokens is key for understanding the semantics of code [11].
Range of important features
Different tasks seem to not only focus on different features, but also focus on a different number of features. While the feature importance for solution domain classification and error detection is concentrated on a small number of features, the feature importance is more spread out for clone detection and code repair. For example, the code repair task requires a transfer of knowledge from overlapping keywords, names, and more generally different tokens from a source language for fixing a code in a transfer language. On the other hand, solution domain classification focuses on fewer features, related to the problems attributes.
Comparison across categories
Comparing the importance of different feature categories across tasks, we observe a relatively higher significance of syntactic features as demonstrated by the top two features Keywords and Names. Compared to transfer learning across natural languages [17], model-specific and linguistics features seem to carry less significance for transfer in programming languages. While de Vries et al. [17] find the potential of cross-lingual transfer dropping significantly for sources and targets that are not seen by the model, our finding brings hope to not only transferring learning to low-resource target languages, but even finetuning the model on unseen languages that could serve as good sources, such as Kotlin.
Within-category differences
While different categories seem to have a higher impact on the model predictions, the impact of single features within a category varies across task. Our results are consistent with prior work that found Names and Keywords to be the most important features for cross-lingual transfer in certain tasks [11]. However, half of the tasks show keywords to be more important than names, while the other half has the opposite order of importance.
1. The importance of language pair features for cross-lingual transfer performance depends on the task. 2. Higher-level tasks seem to draw upon a broader range of features to best understand the code. 3. Overall, the most important features across the four tasks studied in this paper are keywords and names.
V Threats to Validity
V-A External validity
One threat is that to make the large number of experiments feasible and reduce their carbon footprint, we had to limit ourselves to a moderately-sized model. We mitigate this threat by picking CodeT5 [22], which is widely used in the research literature for AI-for-Code. We believe that even if we had used larger models, insights from one model would not perfectly generalize to another.
Another threat is that to cover many languages, we restricted ourselves to four tasks. We mitigate this threat by picking tasks that cover a broad spectrum of difficulty and are diverse, as evidenced by their different performance and feature importances. Overall, our number of tasks exceeds prior work on cross-lingual transfer for programming languages.
A further threat is that our datasets have some synthetic aspects, such as type-IV clones [25] for clone detection data or fault injection for code repair data. However, using fully natural data is such a high standard that hardly any academic paper satisfies it, nor is it typically the case in industry practice. Instead, we mitigate this threat by explicitly measuring the importance of dataset-specific features of language pairs.
Another threat is that we focused on the setting of zero data in the target language. This represents the lowest barrier to entry, as gathering even a small number of examples can be tricky. Therefore, we leave exploring the case where there is some target-language data to future work.
V-B Internal validity
We believe that our study avoids one common internal validity threat, which revolves around insufficient runs. Thanks to the sheer number of language pairs for each task, we get thousands of different experiments, whose results are consistent with each other. Moreover, each of these experiments involves fine-tuning with tens of thousands of samples, giving us confidence that the choice of samples is not a threat either.
Since our study is the first of its kind at this scale, there is a threat that it cannot be directly compared to baselines. We mitigate this by observing that our results do confirm more narrow studies where applicable, for instance, a study of transfer for identifier-centric tasks [11].
VI Conclusion
We perform a systematic and extensive study of LLM transfer learning covering up to 41 programming languages, across 4 tasks including both classification and generation. To understand relative differences in cross-lingual performance between different programming languages, we define several linguistic, syntactic, dataset, and model-specific features of language pairs and then analyze the feature importance for a model that predicts transfer performance. Overall, we hope that this paper helps advance the community’s understanding of how learning transfers among programming languages, and that this improved understanding leads to better models to assist users of those languages, particularly low-resource ones.
VII Data Availability
The experiments are based on the publicly available CodeT5-base (220M parameters) model [22] and the open-sourced datasets CodeNet [19] and XCodeEval [20]. Upon paper acceptance, we plan to open-source the code to create the derived datasets and the framework to train and evaluate the models. Also, we plan to make the 58 finetuned models available with the final version of the paper.
We also plan to submit a replication package to the artifact committee.

### v3

@@
-An Empirical Study of Cross-Lingual Transfer Learning in Programming Languages
+Cross-lingual Transfer in Programming Languages: An Extensive Empirical Study
 Abstract
-Large language models (LLMs) have become remarkably good at improving developer productivity for high-resource programming languages.
-These models use two kinds of data: large amounts of unlabeled code samples for pre-training and relatively smaller amounts of labeled code samples for fine-tuning or in-context learning.
-Unfortunately, many programming languages are low-resource, lacking labeled samples for most tasks and often even lacking unlabeled samples.
-Therefore, users of low-resource languages (e.g., legacy or new languages) miss out on the benefits of LLMs.
-Cross-lingual transfer uses data from a source language to improve model performance on a target language.
-It has been well-studied for natural languages, but has received little attention for programming languages.
-This paper reports extensive experiments on four tasks using a transformer-based LLM and 11 to 41 programming languages to explore the following questions.
-First, how well does cross-lingual transfer work for a given task across different language pairs.
-Second, given a task and target language, how should one choose a source language.
-Third, which characteristics of a language pair are predictive of transfer performance, and how does that depend on the given task.
-Our empirical study with 1,808 experiments reveals practical and scientific insights, such as Kotlin and JavaScript being the most transferable source languages and different tasks relying on substantially different features.
-Overall, we find that learning transfers well across several programming languages.
+Large language models (LLMs) have achieved state-of-the-art performance in various software engineering tasks, including error detection, clone detection, and code translation, primarily leveraging high-resource programming languages like Python and Java.
+However, many critical languages, such as COBOL, as well as emerging languages, such as Rust and Swift, remain low-resource due to limited openly available code.
+This scarcity hampers the training and effectiveness of LLMs for these languages, increasing software maintenance costs and stifling innovation.
+Addressing this gap, we investigate the potential of transfer learning to enhance LLM performance on low-resource programming languages by leveraging data from high-resource counterparts.
+Our extensive empirical study evaluates transferability across 10 to 41 programming languages and five key tasks: code generation, clone detection, code repair, solution domain classification, and error detection.
+Additionally, we develop a performance prediction model to guess the best source languages for a given target and task, and analyze the features that influence transfer performance.
+We further replicate a representative subset of experiments with a larger model to test the generalizability of our conclusions to contemporary large-scale LLMs.
+Our findings demonstrate that cross-lingual transfer significantly outperforms zero-shot learning, with effectiveness varying based on both source and target languages.
+Languages such as Java and Go emerge as the best targets, while Kotlin and JavaScript are excellent sources.
+Furthermore, our model reliably predicts successful transfer sources by considering linguistic and dataset-specific features, offering practical guidance for data acquisition and model training.
+This work contributes to the development of LLM-driven tools for low-resource programming languages and provides insights into the characteristics that facilitate transfer across language pairs.
+1 Introduction
+Large language models (LLMs) leverage the fact that software code is often readable and repetitive (much like natural language (Hindle et al., 2016)) to achieve state-of-the-art performance on many software engineering tasks such as error detection, clone detection, and code translation (Lu et al., 2021).
+However, LLMs have thus far been trained and evaluated mainly on high-resource programming languages such as Python and Java with vast amounts of openly available code.
+Yet there are many low-resource programming languages that lack enough openly available code to train LLMs.
+COBOL, which supports many critical financial applications and may have over 775 billion lines of code overall (MicroFocus, 2022), is low-resource due to lack of open availability.
+Notwithstanding their popularity (Zhou et al., 2022; Coursera, 2024), new languages such as Rust and Swift start out low-resource, by definition.
+Other languages such as R are low-resource because most code in that language has a license that inhibits use for LLM training (Li et al., 2023).
+Besides being underrepresented in training of LLMs, organizations often spend significant resources on training software developers to work with unfamiliar low-resource languages.
+This adds to the cost of software maintenance and may also stifle innovation (Edwards & King, 2021).
+There is a need for state-of-the-art AI tools to enhance developer productivity for low-resource languages.
+Recent works have shown the possibility of transfer learning: leveraging data from one programming language to compensate for the lack of data in another (Ahmed & Devanbu, 2022; Yuan et al., 2022; Pian et al., 2023; Cassano et al., 2024).
+Another recent work has developed several similarity metrics to decide which high-resource programming language dataset can be used to augment fine-tuning data for a target language (Chen et al., 2022).
+Prior work, however, has used at most 6 programming languages, none of which are truly low-resource.
+Further, the notion of similarity among programming languages is underexamined: more research is needed to make reliable claims.
+To overcome these shortcomings, we examine transferability from source to target language combinations for 5 tasks and for many more languages than prior work (10–41 depending on the task).
+We start by using training data for a task in a source language to fine-tune a small LLM (for 4 tasks) or to perform in-context learning using a larger LLM (for the 5th task).
+Then, we feed the respective LLM disjoint test data for the same task on a target language.
+Using this methodology, we evaluate the results across all source languages, target languages, and tasks.
+The source languages are, by definition, high-resource languages since we have enough data to train a model on them.
+Target languages include both high-resource and low-resource ones.
+The finetuning experiments focus on the four tasks clone detection, code repair, solution domain classification, and error detection.
+The in-context learning experiments use a fifth task, code generation.
 Figure 1: Overview.
@@
 For RQ3, measure how important the features of language pairs are.
-Repeat for several languages and four tasks: Error Detection, Solution Domain Classification, Clone Detection, and Code Repair.
-I Introduction
-Large language models (LLMs) leverage the naturalness of software [1] to achieve state-of-the-art performance on many software engineering tasks such as error detection, clone detection, and code translation [2, 3].
-However, LLMs have thus far been trained and evaluated mainly on programming languages with vast amounts of openly available code rather than low-resource ones.
-Programming languages used for critical and strategic applications may have much code only in private repositories and are thus ignored.
-COBOL, which supports many critical financial applications and may have over 775 billion lines of code overall [4], is low-resource due to lack of open availability.
-Similarly with Fortran, which is used in high performance and scientific computing domains.
-Notwithstanding their popularity [5, 6], new languages such as Rust and Swift start out low-resource, by definition.
-Other languages such as R may be low-resource for licensing reasons [7].
-Besides being underrepresented in training of LLMs, organizations often spend significant resources on training software developers to work with unfamiliar low-resource languages.
-This adds to the cost of maintaining such software and may also stifle innovation [8].
-There is a need for state-of-the-art AI tools to augment the capabilities of developers for these low-resource languages.
-Indeed, LLMs are being used in tools to enhance developer productivity and can also help with migration or modernization of projects from one programming language to another by supporting code translation [9, 10].
-Recent works have shown the possibility of transfer learning: leveraging data from one programming language to compensate for the lack of data in the target programming language [11, 12, 13].
-Another recent work has developed several similarity metrics to decide which high-resource programming language dataset can be used to augment fine-tuning data for some other language [14].
-Prior work, however, has used at most 6 programming languages, none of which are truly low-resource.
-Further, the notion of similarity among programming languages is underexamined: more research is needed to make reliable claims.
-To overcome these shortcomings, we examine transferability from source to target language combinations for 4 tasks and for many more languages than prior work (11–41 depending on the task).
-The four tasks we focus on are clone detection, code repair, solution domain classification, and error detection.
-Fig.
-1 gives a high-level overview of the methodology for our extensive empirical study.
-From left to right, we start with a pretrained LLM that has seen unlabeled code in several programming languages.
+Repeat for several languages and four tasks.
+Figure 1 summarizes the methodology for the finetuning part of our empirical study, which builds upon the work of de Vries et al.
+(2022) by exploring programming languages and multiple tasks.
+On the left, we start with a pretrained LLM that has seen unlabeled code in several programming languages.
 Next, for each of four tasks, for each source language with a suitable amount of labeled training data for that task, we finetune the LLM.
 For each target language with a suitable amount of labeled test data for the task, we evaluate the finetuned LLM.
-For example, Fig.
-1 illustrates this with a solution domain classification task, using C as the source language and Rust as the target language, which yields an F1 score of $0.48$.
+For example, Figure 1 illustrates this with a solution domain classification task, using C as the source language and Rust as the target language, which yields an F1 score of $0.48$.
 Doing this for four different tasks and all their source and target language combinations yields four different heat maps.
-Our first research question (RQ1) asks how learning transfers across programming languages in general, by exploring these heat maps.
-Our second research question (RQ2) asks whether one can predict the ranking of source languages in the heat maps from the previous experiment just using features of language pairs, by training a performance prediction model from the language pair features and the ground-truth target labels from the heat maps.
-For example, in Fig.
-1, the performance prediction model for the solution domain classification task might rank C behind Kotlin as a source language for transfer to Rust.
-Given the cost of training LLMs, the performance prediction model is useful in directing data acquisition efforts and deciding how to spend compute resources.
-Finally, as our third research question (RQ3), we seek software-engineering insights into the characteristics of programming languages that affect transferability.
-To do this, we measure the importance of language pair features in the prediction model from the previous experiment.
+The methodology using in-context learning on the fifth task is analogous and yields a fifth heat map.
+This paper addresses three research questions (RQ1–RQ3).
+RQ1: How well does cross-lingual transfer work for a given task across different language pairs? We answer this question directly by exploring the heat maps.
+RQ2: Given a task and target language, how should we pick the source language for best performance?
+We answer this question by training a performance prediction model from the language pair features and the ground-truth target labels from the heat maps, which predicts the ranking of source languages.
+For example, in Figure 1, the performance prediction model for the solution domain classification task might rank C behind Kotlin as a source language for transfer to Rust.
+Given the cost of training an LLM, the performance prediction model is useful in directing data acquisition efforts and deciding how to spend compute resources.
+RQ3: Which characteristics of a language pair are predictive of transfer performance, and how does that depend on the given task?
+We answer this question by measuring the importance of language pair features in the prediction model from the previous experiment.
 This paper makes the following contributions:
-  • Evaluating the pairwise transferability for several programming languages (including low-resource ones) and four tasks using LLMs.
-  • Developing a method to identify the best language to transfer from, for different target languages and tasks.
+  • Evaluating the pairwise transferability for several programming languages (including low-resource ones) and five tasks using language models.
+  • Developing a method to predict the best language to transfer from, for different target languages and tasks.
   • Characterizing the features of programming language pairs that are predictive of transferability for given tasks.
-We are planning to make our data generation, training, and analysis code available.
 One goal of this paper is to offer practical guidance to engineers who build LLM-driven software engineering tools towards more informed choices for data acquisition and modeling.
-A second goal of this paper is to help advance software engineering as a science by using the lens of transfer learning to shed light on programming language characteristics.
-Below are some results highlights:
-  • Learning transfers well for all tasks. More specifically, cross-lingual learning transfers better than zero-shot.
-  • Transfer learning depends on the target language. Java, Go, Dart, and TypeScript are among the best target languages. C++, Python, Vim, and COBOL are among the worst.
+A second goal of this paper is to use the lens of transfer learning to shed light on programming language characteristics.
+Some highlights of our results are as follows:
+  • Learning transfers well for all tasks and cross-lingual learning transfers better than zero-shot.
+  • Transfer learning depends on the target language. Java and Go are the best target languages whereas C++ and Python are the worst.
   • Transfer learning also depends on source language used for finetuning. Kotlin and JavaScript are the best source languages, C++ is the worst.
-  • We can predict which source language can perform well for a given task by considering a variety of features.
-  • Different tasks rely on substantially different features.
-  • Overall the most important features are the overlap of keywords and names between the source and target languages.
-If we take the premise that software engineers will increasingly benefit from LLM support, then we hope this paper will help democratize such support into low-resource settings.
-II Background and Related Work
-Cross-lingual Transfer for Natural Languages
+  • We can reliably predict which source languages will perform well on a given task by considering linguistic, synthetic, dataset-specific, and model-specific features.
+  • Different tasks rely on different features; for some, keywords and names are most important.
+Assuming software engineers increasingly benefit from LLM support, we believe this paper will help democratize such support into low-resource settings and help inspire more research in this area.
+2 Related Work
+Cross-lingual transfer for natural languages.
 Lin et al.
-[15] use 4 tasks to study transfer among up to 60 natural languages and explore how important features of language pairs are for predicting how well learning transfers between them, finding that feature importances vary a lot across tasks.
+(2019) use 4 tasks to study transfer among up to 60 natural languages and explore how important features of language pairs are for predicting how well learning transfers between them, finding that feature importances vary a lot across tasks.
 Lauscher et al.
-[16] use 5 tasks to study transfer among up to 15 languages and argue that while zero-shot performance can work for low-level tasks, higher-level tasks benefit from at least a few target-language shots.
+(2020) use 5 tasks to study transfer among up to 15 languages and argue that while zero-shot performance can work for low-level tasks, higher-level tasks benefit from at least a few target-language shots.
 de Vries et al.
-[17] use 1 task (POS tagging) with 65 source and 105 target languages and study the effects of language families and writing systems, with Romanian emerging as a particularly good source language.
+(2022) use 1 task (POS tagging) with 65 source and 105 target languages and study the effects of language families and writing systems, with Romanian emerging as a particularly good source language.
 Ahuja et al.
-[18] use 11 tasks with 1 source and varying numbers of target languages, confirming earlier findings that feature importances vary a lot across tasks.
-Our paper takes inspiration from these works, especially de Vries et al.
-[17], but differs by studying programming languages instead of natural languages, which have different tasks and different features of language pairs affecting transferability.
-Cross-lingual Transfer for Programming Languages
+(2022) use 11 tasks with 1 source and varying numbers of target languages, confirming earlier findings that feature importances vary a lot across tasks.
+Our paper takes inspiration from these works but differs by studying programming languages instead of natural languages, which have different tasks and different features affecting transferability.
+Cross-lingual transfer for programming languages.
 Zhou et al.
-[5] call for studying transfer, motivating with 1 task (code completion) and 2 languages (from Hack to Rust).
+(2022) call for studying transfer, motivating with 1 task (code completion) and 2 languages (from Hack to Rust).
 Chen et al.
-[14] use 2 tasks (code summarization and search) and 6 languages to study transfer to Ruby, and propose picking the source language based on language similarity.
-Ahmed et al.
-[11] use 3 tasks to study transfer among 6 languages, demonstrating that due to the nature of their tasks, signals from identifiers are highly important for transferability.
+(2022) use 2 tasks (code summarization and search) and 6 languages to study transfer to Ruby, and propose picking the source language based on language similarity.
+Ahmed & Devanbu (2022) use 3 tasks to study transfer among 6 languages, demonstrating that due to the nature of their tasks, signals from identifiers are highly important for transferability.
 Yuan et al.
-[12] use 1 task (automated program repair) to study transfer among 5 languages, sequentially fine-tuning on multiple languages with innovative tricks to prevent catastrophic forgetting.
+(2022) use 1 task (automated program repair) to study transfer among 5 languages, sequentially fine-tuning on multiple languages with innovative tricks to prevent catastrophic forgetting.
 Pian et al.
-[13] use 2 tasks (code summarization and completion) to study transfer among 4 languages, using meta-learning to improve a base learner.
-Our paper also focuses on programming languages, but considers more tasks and vastly more languages to explore conditions for effective transfer.
-Besides the above works, there is also work on learning for multiple programming languages.
-While such prior work has been instrumental in making studies like ours possible, it differs by not focusing on cross-lingual transfer.
-Our work benefits from datasets for multiple programming languages.
-CodeNet covers 55 languages [19]; XCodeEval covers 7 tasks with up to 17 languages [20]; and MultiPL-E supports 1 task (code generation) in 19 languages [21].
-There are also models for multiple programming languages.
-TransCoder uses pre-training across 3 languages to help transpilers [10].
-CodeT5 is pre-trained on 8 languages and 5 tasks [22].
-StarCoder is pretrained on 86 languages [7], but e.g.
-does not include COBOL and evaluation is largely focused on Python.
+(2023) use 2 tasks (code summarization and completion) to study transfer among 4 languages, using meta-learning to improve a base learner.
+Cassano et al.
+(2024) introduce MultiPL-T, a new semi-synthetic training dataset, and show that it improves transfer for 1 task (NL to code completion) across 19 languages.
+Our paper also focuses on programming languages, but considers more tasks and more languages to explore conditions for effective transfer.
+There is also work on learning across multiple programming languages that has been instrumental in making studies like ours possible, but it differs in that it does not focus on cross-lingual transfer.
 Unsupervised pretraining on multiple languages has become common, but transferability of supervised tasks has not yet been thoroughly studied; our paper addresses that gap.
-III Experimental Setup
-As illustrated in Fig.
-1, given a task, our experimental approach first finetunes a model for each source language individually and then tests each finetuned model on all target languages.
+One could also imagine exploring transfer from multiple source languages to a low-resource target; we leave this to future work.
+3 Experimental Setup
+Most of our experiments are based on finetuning a small model for four tasks as illustrated in Figure 1; in addition, there are also some experiments based on few-shot prompting a larger model without finetuning.
+For most experiments, given a task, our experimental approach first finetunes the CodeT5 (Wang et al., 2021) model for each source language individually and then tests each finetuned model on all target languages.
 We applied this approach to four tasks where the number of source languages varies from 6 to 22, leading to 58 finetuned models in total.
-For each source language, we finetune the model using one A100 GPU for 6-30 hrs depending on the task.
+For each source language, we finetune the model using one A100 GPU for 6 to 30 hours depending on the task.
 Each model was then evaluated on 11 to 43 target languages producing 1,808 experiments.
-All results are presented in Figs.
-5(a)–(d) and analyzed in Sec.
-IV.
-This section presents methodological details.
+All results are presented in Figures 5(a)–(d) and analyzed in Section 4.
+For the few-shot prompting experiments, we generated responses with the Llama 3.3 70B-Instruct model via TogetherAI, 2025 using a temperature of 0.8.
+Those results are presented in Figure 6.
 (a) Clone Detection: 21 source languages $\times$ 41 target languages. Metric: F1 Score.
@@
 The figure shows scores for every combination of source and target language.
-Each column corresponds to a source language, with “zsh” showing zero-shot performance, i.e., without fine-tuning on any source language.
+Each column corresponds to a source language, with “Zero Shot” showing zero shot performance, i.e., without fine-tuning on any source language.
 Each row corresponds to a target language.
@@
 (a) Code Repair: 20 source languages $\times$ 38 target languages. Metric: BLEU score.
-Figure 3: Transfer scores heat map (cont.; see caption from 2(a))
+Figure 3: (continued; see caption from 2(a))
 (a) Solution Domain Classification: 11 source languages $\times$ 11 target languages. Metric: F1 score.
-Figure 4: Transfer scores heat map (cont.; see caption from 2(a)).
+Figure 4: (continued; see caption from 2(a))
 (a) Error Detection: 6 source languages $\times$ 11 target languages. Metric: F1 Score.
-Figure 5: Transfer scores heat map (cont.; see caption from 2(a)).
-Base Datasets
-The four tasks studied are derived from two multilingual code datasets: CodeNet [19] and XCodeEval [20].
-CodeNet [19] consists of about 14M code samples for a total of 500M lines of code in 55 different programming languages, derived from submissions to online coding judge websites.
+Figure 5: (continued; see caption from 2(a))
+(a) Code Generation: 10 source languages $\times$ 10 target languages. Metric: Pass@1 Score.
+Figure 6: Few shot prompting transfer scores heat map.
+The figure shows scores for every combination of source and target language.
+Each column corresponds to a source language, with “Zero Shot” showing zero shot performance.
+Each row corresponds to a target language.
+Similarly, dendrograms show results of hierarchical clustering and framed black boxes highlight the performance of a source language on itself as the target language.
+Base datasets.
+The five tasks studied are derived from three multilingual code datasets: CodeNet, XCodeEval, and MBXP.
+CodeNet (Puri et al., 2021) consists of about 14M code samples in 55 different programming languages, derived from submissions to online coding judge websites.
 The dataset comes with benchmarks for code classification and code similarity.
-XCodeEval [20] contains 25M samples in 11 programming languages from another different online coding judge website.
+XCodeEval (Khan et al., 2023) contains 25M samples in 11 programming languages from another different online coding judge website.
 This dataset comes with an execution-based evaluation framework and several different benchmarks: from classification (solution domain classification and error detection) to generation (program synthesis, automatic program repair, and code translation).
-Tasks
-To study how learning transfers between programming languages, we explore two type of tasks: classification (error detection, solution domain classification, clone detection) and generation (code repair).
+MBXP (Athiwaratkun et al., 2023) is a multilingual extension of MBPP (Austin et al., 2021) to Java, JavaScript, TypeScript, Go, Ruby, Kotlin, PHP, C#, Scala, C++, Swift, and Perl.
+A companion code package is provided to perform execution in all supported languages.
+The dataset is designed for investigating the effectiveness of few-shot prompting, zero-shot translation, as well as the robustness to prompt perturbation.
+In our experiment, we utilize the few shot prompting setup.
+Tasks.
+To study how learning transfers between programming languages, we explore two type of tasks: classification (error detection, solution domain classification, clone detection) and generation (code repair, code generation).
 For the classification tasks, we follow Lewis et al.
-[23] by predicting labels from the vocabulary of class labels based on the final decoder hidden state.
-We use BLEU score [^1] We chose BLEU score because it enables experiments with the larger number of languages in CodeNet [19], and because other prominent benchmarks for the code repair task also adopt this metric [24].
-for evaluating the generative task and F1 score for evaluating the classification tasks.
-  • Error Detection: Given a code $C$ in language $L$, the task is binary classification of whether or not $C$ compiles (or can be loaded by an interpreter) without error.
-  • Solution Domain Classification: Given a code $C$ in language $L$, the task is multi-label classification of $C$ into a set of tags corresponding to potential algorithmic techniques required to write the solution (e.g., sorting, graphs).
-  • Clone Detection: Given two code samples $C_{1}$ and $C_{2}$ in language $L$, the task is binary classification of whether they are type-IV clones (semantically similar) [25].
-To generate the dataset, we apply the following procedure to the CodeNet dataset.
+(2020) by predicting class labels based on the final decoder hidden state.
+We use F1 score for evaluating the classification tasks, which is a widely accepted practice since it is good at dealing with class imbalance.
+We use BLEU score for evaluating the code repair task; while an execution-based metric would provide value, it was infeasible due to the large number of languages in CodeNet (Puri et al., 2021).
+(Other prominent benchmarks for the code repair task also adopt BLEU score (Lu et al., 2021).) Finally, we use execution-based Pass@1 score for evaluating the code generation task.
+The execution-based evaluation checks the functional correctness of the generated code by running it against the provided tests.
+Error Detection: Given a code $C$ in language $L$, do binary classification of whether $C$ compiles (or can be loaded by an interpreter) without error.
+Solution Domain Classification: Given a code $C$ in language $L$, do multi-label classification into a set of tags corresponding to algorithmic techniques required to write the solution (e.g., sorting, graphs).
+Clone Detection: Given two code samples $C_{1}$ and $C_{2}$ in language $L$, do binary classification of whether they are type-IV clones (semantically similar) (Roy & Cordy, 2007).
+We derived the dataset from CodeNet as follows.
 Given all combinations of solutions to all problems in language $L$, we identify positive samples (clones) as pairs of accepted solutions for the same problem and the others as negative examples.
-To balance the positive samples across problems, we ensure a ratio of 0.15 of positive samples across different languages.
-  • Code Repair: Given a buggy code $C$ in language $L$, the task is to generate the corresponding repaired code.
-We generate the dataset by using the following procedure on CodeNet solutions: Each sample is modified by sequentially inserting, removing, or replacing tokens of different types for fixed ratios for different token types.
-Data Sampling
+To balance the data across problems, we ensure a ratio of 0.15 of positive samples across different languages.
+Code Repair: Given a buggy code $C$ in language $L$, generate the corresponding repaired code.
+We derived the dataset from CodeNet, modifying samples by sequentially inserting, removing, or replacing tokens of different types for fixed ratios for different token types.
+Code Generation: Given a prompt $P$ consisting of a function signature and a docstring in language $L$, the task is to generate the function body completion.
+In the few shot prompting setup, few-shot prompts consisting of three correct functions from the respective MBXP dataset are provided.
+Few shot examples precede the function completion prompt for each evaluation.
+Data sampling.
+Our objective is to facilitate an extensive study on transfer across the maximum number of language pairs, which can be helpful for those working with low-resource languages.
+Sources are selected based on the relative performance of fine-tuned checkpoints compared to a random baseline, in order to determine if there is enough signal in the data to facilitate learning.
+Languages that demonstrate higher performance than the baseline are included as sources.
 Given variation in the size of datasets for different languages, we follow the sampling procedure of de Vries et al.
-[17] to select the number of training samples and identify potential source languages.
+(2022) to select the number of training samples and identify potential source or high-resource languages.
 We first finetune the model with datasets of different number of samples $N$={10K, 30K, 50K, 70K, 100K}.
-Depending on the number of training examples $N_{L}$ for a language $L$, we randomly upsample languages with $N<N_{L}$ or downsample languages with $N>N_{L}$.
-We filter the set of source languages for a given task based on the relative performance compared to a baseline model.
+Depending on the number of training examples $N_{L}$ for a language $L$, we randomly upsample languages with We filter the set of source languages for a given task based on the relative performance compared to a baseline model.
 We select sample sizes of 50K for solution domain classification, 70K for error detection, and 100K for clone detection and code repair.
-Fig.
-8 in the supplementary material shows how the metric for each task vaires with the sample size.
-Model
-Our experimental framework is based on CodeT5-base (220M parameters) leveraging the Hugging Face transformers library [26].
+Models.
+Our finetuning experiments are based on CodeT5-base (220M parameters) using the Hugging Face transformers library (Wolf et al., 2020).
 Code-T5 is an open-source model, pre-trained on eight programming languages (Ruby, JavaScript, Go, Python, Java, PHP, C, and C#).
-Due to its encoder-decoder nature, the model performs well on both code generation and code understanding tasks [22] like code repair, error detection, and clone detection, among others.
-This, along with its relatively small size, makes it a good fit for our empirical study, which requires 58 finetuned models with numerous inference runs each on four tasks.
-IV Results and Discussion
-This section addresses the following research questions.
-RQ1:
-How well does cross-lingual transfer work for a given task across different language pairs?
-RQ2:
-Given a task and target language, how should we pick the source language for best performance?
-RQ3:
-Which characteristics of a language pair are predictive of transfer performance, and how does that depend on the given task?
-IV-A Transfer analysis
+We consider these languages as seen languages and Figure 5 highlights the names of these pre-training languages in red font.
+Due to its encoder-decoder nature, the model performs well on both code generation and code understanding tasks (Wang et al., 2021) like code repair, error detection, and clone detection, among others.
+This, along with its relatively small size, makes it a good fit for our empirical study, which requires 58 finetuned models with numerous inference runs each.
+We keep the same hyperparameters for all the experiments: learning rate of 2e-5, batch size of 8, and the number of epochs set to 20.
+The few-shot prompting experiment use Llama 3.3 70B-Instruct, [^1] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct an instruction-tuned multilingual LLM (70B parameters) (Grattafiori et al., 2024).
+Data Splits
+Language Train Test
+php 914,924 1,975
+rb 912,159 3,963
+hs 843,011 2,045
+py 785,760 4,235
+kt 731,575 1,552
+js 713,854 1,600
+c 555,924 2,363
+scala 473,594 1,705
+go 352,181 867
+java 266,248 1,338
+d 247,954 1,355
+cs 233,059 1,151
+pl 185,685 1,762
+nim 175,839 765
+Language Train Test
+ml 175,056 2,008
+f 172,046 1,253
+cpp 150,492 1,792
+pas 136,961 1,370
+rs 128,271 478
+awk 95,312 1,035
+jl 88,700 1,737
+sh 64,899 1,070
+ts 47,414 799
+l 45,986 585
+swift 38,531 737
+fs 31,001 962
+scm 22,787 826
+cr 21,733 1,356
+Language Train Test
+sed 21,584 1,349
+lua 20,688 1,253
+pyx 18,184 931
+dc 17,322 1,706
+vb 7,003 1,132
+octave 5,370 1,476
+vim 4,176 745
+cob 3,437 1,038
+clj 1,058 910
+dart 506 480
+moon 242 453
+ex 330 549
+Table 1: Number of train and test samples for Clone Detection for each language.
+Train and test splits for Error Detection and Solution Domain Classification are provided from the original dataset as described in the dataset statistics from xCodeEval (Khan et al., 2023).
+For Code Repair, 50,000 training examples and 1,000 test examples are synthetically generated for each language.
+For Clone Detection, a distinct set of problems is used for train and test splits, including languages with a minimum of 450 test examples.
+Table 1 shows the number of examples for Clone Detection for each language.
+In the few-shot prompting experiment, we conduct a cross-lingual evaluation across all language pairs, using nearly 1,000 examples for each.
+We additionally use the provided synthetic canonical solutions to measure dataset features across overlapping problem solutions.
+4 Results and Discussion
+This section addresses the research questions RQ1–RQ3 stated in the introduction.
+4.1 Transfer analysis
 We explore RQ1 via extensive experiments on every combination of source and target programming language, for each task.
-The four heat maps of Fig.
-5 show the results.
+The four heat maps of Figure 5 show the results.
 A source programming language is the language used for fine-tuning a model, with samples in the training data.
 A target programming language is the language used to evaluate a model, with samples in the test data.
-Fig.
-5 shows source languages on the X-axis and target languages on the Y-axis.
+Figure 5 shows source languages on the horizontal axis and target languages on the vertical axis.
 Monolingual scores (where the source and the target language are the same) are highlighted with bold boxes.
+The first column from the left indicates zero-shot performance of the base Code-T5 model on different target languages.
 For RQ1 we analyze model performance trends with respect to task, source language, and target language.
-We provide some practical insights, like which languages are the toughest target languages and which are the best source languages.
-We also discuss some interesting source-and-target language pairs.
-Task dependency
-The heat maps in Fig.
-5 show that transferability of source languages varies depending on the task.
+Task dependency.
+The heat maps in Figure 5 show that transferability of source languages varies depending on the task.
 Visually this can be seen by how the color gradient varies for each task.
@@
 The mean cross-lingual scores are 0.78 for clone detection, 0.75 for code repair, 0.65 for error detection, and 0.47 for solution domain classification.
-The monolingual scores (in bold boxes) are higher (more red) than the cross-lingual scores, averaging 0.91 for clone detection, 0.98 for code repair, 0.78 for error detection, and 0.51 for solution domain classification.
-The zero-shot columns, where the model is not finetuned on any source language, have lower scores (more blue) for all heat maps.
-The mean zero-shot scores are 0.49 for clone detection, 0.28 for code repair, 0.49 for error detection, and 0.01 for solution domain classification.
+The monolingual scores (in bold boxes) are higher (more red) than the cross-lingual scores, averaging 0.91, 0.98, 0.78, and 0.51 for the same four tasks.
+The zero-shot columns, where the model is not finetuned on any source language, have lower scores (more blue) for all heat maps, averaging 0.49, 0.28, 0.49, and 0.01.
 Variations in the mean monolingual, cross-lingual, and zero-shot scores across tasks confirm that model performance is task dependent.
-The difference in overall zero-shot and cross-lingual scores indicates that in the absence of fine-tuning data for any task in any programming language, it is better to finetune with some other language in that task than using a zero-shot setting.
-This cross-lingual advantage over the zero-shot setting occurs for every task.
+The difference in overall zero-shot and cross-lingual scores indicates that in the absence of fine-tuning data for any task in any programming language, it is better to finetune with some other language in that task than to use a zero-shot setting.
+This cross-lingual advantage holds for every task.
+Mean Score Task Mono Cross Overall Zero-Shot Clone Detection 0.91 0.78 0.79 0.49 Code Repair 0.98 0.75 0.76 0.28 Error Detection 0.78 0.65 0.66 0.49 Solution Domain Classification 0.51 0.47 0.47 0.01
+Table 2: Mean Scores by Task.
+Monolingual implies finetuning where the train and test data language is the same.
+Cross-lingual implies finetuning where the train and test data languages differ.
+Overall scores include both monolingual and cross-lingual scenarios.
+Zero-shot means the performance of the base pre-trained model (CodeT5-220M) on the test set without finetuning.
 Looking at these scores in absolute terms reveals that the choice of tasks also affects transferability.
@@
 Within each task, model performance depends on the choice of source and target programming language, which we explore next.
-Tab.
-II in the supplementary material shows the mean score by task.
-Target language dependency
-Target programming languages fall into two categories, high- and low-resource.
-In Fig.
-5, high-resource target programming languages are represented on both axes.
-Low-resource target programming languages are represented only on the Y-axis.
+Table 2 shows the mean score by task.
+Learning transfers well for all tasks. Cross-lingual learning works better than zero-shot.
+Target language dependency.
+Clone Detection Code Repair Error Detection Solution Domain Target cross 0-shot cross 0-shot cross 0-shot cross 0-shot Mean Rank Java 0.86 0.44 0.84 0.25 0.66 0.60 0.42 0.009 4.25 Go 0.86 0.50 0.71 0.16 0.66 0.55 0.51 0.003 4.25 Rust 0.80 0.49 0.77 0.23 0.66 0.36 0.53 0.001 4.50 JavaScript 0.83 0.48 0.85 0.34 0.65 0.49 0.52 0.008 4.75 Kotlin 0.84 0.49 0.76 0.17 0.66 0.48 0.50 0.005 5.00 PHP 0.84 0.51 0.82 0.28 0.65 0.38 0.50 0.004 5.25 C# 0.84 0.48 0.82 0.2 0.64 0.54 0.47 0.046 5.50 C 0.83 0.45 0.82 0.34 0.66 0.57 0.48 0.004 5.75 C++ 0.79 0.46 0.81 0.24 0.66 0.51 0.41 0.004 7.00 Python 0.83 0.50 0.76 0.20 0.64 0.49 0.37 0.006 8.75
+Table 3: Score distribution by target languages common for all tasks.
+Scores are the mean of the score of every source for a given target language.
+Mean rank is used to rank the languages based on their ranking for each task.
+All the languages shown are high-resource target languages for Clone Detection, Code Repair and Solution Domain Classification, which means they are also among the source languages for those tasks.
+C#, Go, PHP, Ruby, and Rust are low-resource languages for Error Detection.
+The table shows that all languages benefit from transfer learning.
+Java and Go seem to benefit the most and C++ and Python seem to benefit the least.
+In Figure 5, high-resource target programming languages are represented on both axes.
+Low-resource target programming languages are represented only on the vertical axis.
 Models are finetuned only on high-resource programming languages and tested on both high- and low-resource programming languages.
@@
 Four (Go, Rust, PHP, and C#) are low-resource in error detection.
-For each of these ten languages, for each of the four tasks, we calculate the mean cross-lingual score using all scores in the row of Fig.
-5 for each language, excluding the monolingual score in the bold box and the zero-shot score in the extreme left.
-We exclude monolingual and zero-shot scores so as to understand transferability from different languages to the target language.
+For each of these ten languages, for each of the four tasks, we calculate the mean cross-lingual score using all scores in the row of Figure 5 for each language, excluding the monolingual score in the bold box and the zero-shot score in the extreme left.
 Based on this mean we rank each of the 10 languages for each task, and then calculate the mean rank for each language.
-We find Java, Go, Rust, JavaScript, and Kotlin are target languages that benefit most from transfer learning, while PHP, C#, C, C++, and Python benefit least, with Java, Go being the best and Python, C++ being the worst.
-C++ has a reputation as being one of the hardest programming languages to learn for human software developers, which may explain our finding that it is a difficult target for LLM transfer learning.
-Some target programming languages are good at monolingual performance, but much worse at cross-lingual performance; one example across multiple tasks is C++.
-In-language performance for C++ is 0.95, 0.99, and 0.46 for clone detection, code repair, and solution domain classification, whereas its mean cross-language performance is 0.79, 0.81, and 0.41 for the same tasks.
-Java is an example of more consistent in-language and cross-language transferability across tasks and hence can be considered relatively more transferable than C++.
-Java has in-language scores of 0.95, 0.99, and 0.45 for clone detection, code repair, and solution domain classification, and cross-language scores of 0.86, 0.84, and 0.42 for the same.
-All numbers are reported Tab.
-III in the supplementary material.
+The most transferrable target languages from best to worst are Java, Go, Rust, JavaScript, Kotlin, PHP, C#, C, C++, and Python.
+Some target programming languages are good at monolingual performance, but much worse at cross-lingual performance.
+For example, in-language performance for C++ is 0.95, 0.99, and 0.46 for clone detection, code repair, and solution domain classification, whereas its mean cross-language performance is 0.79, 0.81, and 0.41 for the same tasks.
+This might be explained by the fact that C++ has a reputation as being one of the hardest programming languages to learn for human software developers.
+All numbers are in Table 3.
+Clone Detection Code Repair Target Language cross 0-shot cross 0-shot Mean Rank Dart 0.87 0.51 0.79 0.28 2.0 TypeScript 0.83 0.49 0.83 0.25 3.5 Elixir 0.83 0.51 0.75 0.21 5.0 Lua 0.82 0.51 0.77 0.20 5.0 Swift 0.85 0.52 0.72 0.23 6.5 Visual Basic 0.80 0.49 0.75 0.09 7.5 Cython 0.79 0.51 0.75 0.23 7.5 Julia 0.83 0.49 0.72 0.29 8.0 Moonscript 0.74 0.50 0.75 0.41 9.0 Sed 0.69 0.53 0.85 0.60 9.5 Scheme 0.78 0.49 0.71 0.14 11.5 Shell 0.74 0.49 0.74 0.39 12.0 Clojure 0.80 0.48 0.67 0.08 12.5 COBOL 0.73 0.49 0.74 0.03 12.5 F# 0.78 0.49 0.71 0.06 13.0 Octave 0.76 0.50 0.71 0.55 13.0 bc 0.73 0.50 0.67 0.55 16.5 Vim 0.69 0.49 0.67 0.48 16.5
+Table 4: Score distribution for low resource target languages common for Clone Detection and Code Repair.
+Scores are the mean of the score of every source language for the given target language.
+Mean rank is used to rank the languages based on their ranking across each task.
+All languages benefit from transfer learning.
+Dart, TypeScript benefit the most and bc, Vim benefit the least.
 We do a similar analysis for low-resource target languages in clone detection and code repair, since these two tasks have a wider and more varied set of low-resource languages.
-Here we find Dart, TypeScript, Elixir, Lua, and Swift benefit most from cross-lingual training, with Dart being the best.
-Dart, TypeScript have average scores of 0.87 and 0.79, 0.82 and 0.83 for clone detection and code repair.
-COBOL, F#, Octave, bc, and Vim are least likely to benefit, with Vim being the worst.
-Vim has average scores of 0.69, 0.67 and COBOL has average scores of 0.72, 0.74 for Clone Detection and Code Repair.
-Tab.
-IV in the supplementary material has more information on mean rank and average score of low-resource target languages for clone detection and code repair.
+Here, the most transferrable target languages from best to worst are Dart, TypeScript, Elixir, Lua, Swift, COBOL, F#, Octave, bc, and Vim.
+See Table 4 for details.
 For error detection, only a smaller set of languages can be considered low resource.
-We observe low variation in average scores across different sources with highest being Rust with 0.66 and lowest being C# with 0.64.
+Average scores vary less across different sources with highest being Rust with 0.66 and lowest being C# with 0.64.
 For solution domain classification there are no low-resource target languages.
-Our analysis shows that transferability depends on the choice of task and target language.
-We can see by comparing with zero-shot scores that all target programming languages benefit from cross-lingual training, however some languages benefit much more than others.
-Most Transferable Source Language
+While all target languages benefit from cross-lingual training, some benefit much more than others. Among low-resource target languages, Dart and TypeScript benefit the most.
+Most Transferable Source Language.
+Clone Detection Code Repair Solution Domain Error Detection Source score rank score rank score rank score rank Mean Rank Kotlin 0.8 3 0.78 3 0.5 1 0.67 1 2.0 JavaScript 0.82 1 0.81 1 0.47 4 0.62 6 3.0 Java 0.8 4 0.78 5 0.49 3 0.66 2 3.5 Python 0.73 6 0.79 2 0.49 2 0.66 5 3.75 C 0.8 2 0.78 4 0.45 5 0.66 4 3.75 C++ 0.79 5 0.78 6 0.44 6 0.66 3 5
+Table 5: Source languages common for all tasks ranked by mean score across target languages and tasks.
+Score is the mean of the scores of all target languages for a given source language.
+We see that Kotlin is relatively the best source language and C++ the worst.
+This result is surprising since Kotlin is not a seen language during pre-training.
 The number of source languages vary for each of the four tasks.
-To understand the source language dependency and to identify the most transferrable source language, we start from languages all four tasks have in common (JavaScript, C, Kotlin, Java, C++, and Python).
+To understand the source language dependency and to identify the most transferrable source language, we start from languages all four tasks have in common (C, C++, Java, JavaScript, Kotlin, and Python).
 For each task, for each language, we calulate the mean cross-lingual score.
-As an example, consider the column for source language Python on the X-axis of Fig.
-5(a).
-We calculate the mean over the scores except 0.94, which is the monolingual score with target language Python.
 Once we have the cross-lingual mean for each language, we rank the source languages based on their mean score for each task.
@@
 The most transferrable source languages from best to worst are Kotlin, JavaScript, Java, Python, C, and C++.
-It may be surprising that Kotlin is the best-performing source language, since it is rarely discussed in the context of training LLMs and is not part of the pre-training corpus for CodeT5 [22].
-Prior work on transfer learning for natural languages [17] yielded a similar surprise, with Romanian as the best-performing source language.
+It may be surprising that Kotlin is the best-performing source language, since it is rarely discussed in the context of training LLMs and is not part of the pre-training corpus for CodeT5 (Wang et al., 2021).
+Prior work on transfer learning for natural languages (de Vries et al., 2022) yielded a similar surprise, with Romanian as the best-performing source language.
 A possible explanation could be that Romanian and Kotlin both have strong roots in dominant languages of the past (Latin and Java), while sitting at a cross-roads that helped them absorb many influences.
-Detailed scores and rankings are in Tab.
-V in the supplementary material.
+Detailed scores and rankings are in Table 5.
+Clone Detection Code Repair Error Detection
+Source score rank score rank score rank Mean Rank
+JavaScript 0.81 1 0.82 1 0.63 6 2.67
+Java 0.79 2 0.75 5 0.66 1 2.67
+Kotlin 0.78 3 0.77 3 0.66 4 3.33
+C 0.77 4 0.76 4 0.66 2 3.33
+Python 0.71 6 0.79 2 0.65 5 4.33
+C++ 0.76 5 0.75 6 0.66 3 4.67
+Table 6: Source languages common for all tasks ranked by mean score across low-resource target languages and three tasks.
+Solution Domain Classification is missing because all the target languages for this task are high-resource.
+Score is the mean of the scores of all target languages for a given source language.
+We see that JavaScript, Java are relatively the best source languages and C++ the worst.
 The above analysis considers all target languages, including high-resource target languages.
@@
 If we repeat the above analysis with only low-resource languages, we are left with a smaller set of target languages and only three tasks (all target languages for solution domain classification are high-resource).
-For the remaining three tasks and six common source languages, analysis shows that the most transferrable source languages are JavaScript, Java, Kotlin, C, Python, and C++, with JavaScript being the best and C++ being the worst.
-JavaScript and Java contribute the most data to CodeT5 pretraining; Kotlin and C++ are absent [22].
-For details see Tab.
-VI in the supplementary material.
+For the remaining three tasks and six common source languages, the most transferrable source languages from best to worst are JavaScript, Java, Kotlin, C, Python, and C++.
+JavaScript and Java contribute the most data to CodeT5 pretraining; Kotlin and C++ are absent (Wang et al., 2021).
+For details see Table 6.
+Clone Detection Code Repair
+Source score rank score rank Mean Rank
+JavaScript 0.81 2 0.82 1 1.5
+Scala 0.81 1 0.77 6 3.5
+Haskell 0.81 4 0.78 3 3.5
+Kotlin 0.78 11 0.77 5 8
+D 0.81 3 0.71 15 9
+Java 0.79 9 0.75 9 9
+PHP 0.79 7 0.74 13 10
+Nim 0.80 5 0.71 16 10.5
+Python 0.71 19 0.79 2 10.5
+Rust 0.79 8 0.72 14 11
+Perl 0.79 6 0.68 17 11.5
+C 0.76 15 0.76 8 11.5
+Go 0.76 16 0.77 7 11.5
+C# 0.77 12 0.74 12 12
+bf 0.55 20 0.77 4 12
+Fortran 0.78 10 0.67 18 14
+Awk 0.76 17 0.75 11 14
+C++ 0.76 18 0.75 10 14
+Lisp 0.77 13 0.67 19 16
+Pascal 0.77 14 0.66 20 17
+Table 7: Source languages common for Clone Detection and Code Repair ranked by mean score across low-resource target languages.
+Score is the mean of the scores of all target languages for a given source language.
+We see that JavaScript, Scala, and Haskell are relatively the best source languages and C++, Lisp, and Pascal the worst.
 If we consider only clone detection and code repair, we get a much broader set of about 20 common source languages.
 Repeating the analysis for these two tasks for low-resource target languages only, we find once again that JavaScript is the most transferable and C++ is among the least.
-In this case the most transferable languages are JavaScript, Scala, and Haskell, while the least transferable are Fortran, AWK, C++, Lisp, and Pascal.
-More details are in Tab.
-VII in the supplementary material.
+More details are in Table 7.
+Kotlin is the best source language over all target languages and tasks. JavaScript is the best source language for low-resource target languages. C++ is a poor source language.
 Language Pair Performance
+Clone Detection Code Repair Error Detection
+Target Source score Source score Source score
+Crystal Ruby 0.93
+Dart Scala 0.93 D 0.88
+Swift JavaScript 0.91 Kotlin 0.79
+bc Scala 0.91 JavaScript 0.87
+TypeScript JavaScript 0.90 JavaScript 0.95
+Elixer C 0.89 JavaScript 0.83
+Clojure L 0.89 bf 0.77
+Visual Basic Scala 0.89 JavaScript 0.81
+Julia Ruby 0.89 JavaScript 0.81
+Cython Python 0.88 Python 0.95
+Lua Ruby 0.88 JavaScript 0.87
+Moonscript Scala 0.87 JavaScript 0.87
+Scheme Lisp 0.86 L 0.83
+COBOL Haskell 0.85 Python 0.78
+F# Haskell 0.84 Haskell 0.80
+Sed perl 0.81 C 0.89
+Octave ml 0.81 Python 0.78
+Bash perl 0.80 bf 0.83
+dc perl 0.80
+Awk Vim 0.78
+C# C++ 0.68
+Rust Python 0.67
+Ruby Kotlin, Java 0.67
+PHP Kotlin, C++, C 0.67
+Go Kotlin 0.67
+Table 8: The best performing source language for each low resource target language, in each task.
+For high resource languages, the best performing source language is usually the same language.
+Ruby is high-resource for Clone Detection and not a target for Code Repair, hence its results are not show for these two tasks.
+Similarly C#, Rust, PHP, and Go are high-resource for Clone Detection and Code Repair.
 The best model performance for each task is almost always in the monolingual setting.
@@
 Another reason to consider individual language pairs is that from the perspective of developers working with low-resource languages, it would be helpful to know which programming languages act as the best source language.
-For clone detection in Fig.
-5(a), notice that a model finetuned on Ruby does almost as well on Crystal, a language whose syntax is inspired by Ruby.
-Fig.
-5(a-b) also shows that surprisingly, Haskell does very well on COBOL, much better than FORTRAN does on COBOL.
-From the same heat maps notice Lisp does very well on Scheme, which can be expected as Scheme is a dialect of Lisp.
+For clone detection in Figure 5(a), a model finetuned on Ruby does almost as well on Crystal, a language whose syntax is inspired by Ruby.
+Figure 5(a-b) also shows that surprisingly, Haskell does very well on COBOL, much better than FORTRAN does on COBOL.
+Lisp does very well on Scheme, which can be expected as Scheme is a dialect of Lisp.
 Similary Python does very well on Cython.
-Fig.
-5(a-d) also show that C-category languages (C, C++, and C#) do well on each other.
-Details are in Tab.
-VIII in the supplementary material.
-Dendrograms
+Figure 5(a-d) also show that C-category languages (C, C++, and C#) do well on each other.
+Details are in Table 8.
+Some language pairs show particularly good results. We need to analyze their features to understand why, motivating Section 4.3.
+Dendrograms.
 Dendrograms for each task cluster programming languages with similar transfer learning performances.
 In the code repair task, syntaxically and semantically related languages are next to each other.
-For example, C is paired with C++ and is next to the pair C# and Java.
-JavaScript is paired with TypeScript, Python with Cython, Lua with Moon, Lisp with Scheme.
+For example, C is paired with C++, JavaScript with TypeScript, Python with Cython, Lisp with Scheme.
 On the other hand, relatedness of languages plays a smaller role in the dendrograms for other tasks.
-This might be because while the code repair task requires deep language understanding, other tasks depend more upon other features.
-For instance, Ahmed and Devanbu [11] have shown that for some tasks, identifiers are more important.
-Section IV-C will analyze the relationship between features of language pairs and their transfer performance.
-1.
-Learning transfers well for all tasks.
-Cross-lingual learning transfers better than zero-shot.
-2.
-Learning depends on the target language.
-Among low resource languages, Dart and TypeScript are the best while Vim and COBOL are among the worst.
-3.
-Kotlin and JavaScript are among the best target languages, C++ is among the worst.
-4.
-Some cross-lingual language pairs show particularly good results, such as Ruby and Crystal, Haskell and COBOL.
-IV-B Performance Prediction
+This might be because while code repair requires deep language understanding, other tasks depend more upon other features.
+For instance, Ahmed & Devanbu (2022) show that some tasks depend heavily on identifiers.
+Relatedness of languages plays a big role for code repair but less so for other tasks.
+4.2 Performance Prediction
+Given a task and target language, how should we pick the source language for best performance?
+Knowing this without having to train models on different source languages can save significant resources.
+It also provides a basis to study important features that characterize a successful transfer from a language to another.
+To predict the performance of the models on the different tasks for different source languages, we train a ranking model based on gradient-boosted decision trees (Ke et al., 2017).
+We consider the features defined in Table 9, grouped into four categories.
+(1) Linguistic features: general properties characterizing a language pair like whether they support the same paradigms, the same style of memory management (garbage collector or not), or the same kind of type-system.
+(2) Syntactic features: properties of the syntax of a language pair measured by overlap in the counts of certain token types, as determined using the Pygments library on the corpus of code in the given languages.
+(3) Dataset-specific features: properties of the problems associated with code samples for the language pair in the online coding judge website from which the dataset came.
+(4) Model-specific features: whether the model saw the source, the target, or both languages during pretraining.
+The linguistic features of each language are given Table LABEL:tab:linguistic-features.
 Feature Measure Description Linguistic Features Object oriented e Is the language object oriented?
@@
 Syntactic Features Name o Number of names which overlap Text o Number of text data which overlap Keyword o Number of keywords which overlap Literal o Number of literals which overlap Punctuation o Number of punctuation signs which overlap Operator o Number of operators which overlap Comment o Number of comment tokens which overlap Syntax o Number of AST nodes which overlap Tokens o Number of tokens which overlap Dataset-Specific Features Difficulty $x$ Average difficulty of dataset problems Length $x$ Average number of tokens Time limit $x$ Average time limit of dataset problems Memory limit $x$ Average memory limit of dataset problems Model-Specific Features Pretrained s Source language is included during pretraining Pretrained t Target language is included during pretraining Pretrained b Both source and target languages in pretraining
-TABLE I: Features of language pairs.
+Table 9: Features of language pairs.
 Column “Measure” indicates how the feature is computed: (e)quivalence, (o)verlap, (s)ource, (t)arget, (b)oth source and target, (rd) for relative difference.
 For dataset specific features, ($x$) can be s, t, or rd.
-Details of the linguistic features are available in supplementary material Tab.
-IX
-Given a task and target language, how should we pick the source language for best performance?
-Knowing this without having to train many models on different source languages can save significant resources.
-It also provides a basis to study important features that characterize a successful transfer from a language to another.
-To predict the performance of the models on the different tasks for different source languages, we train a ranking model based on gradient-boosted decision trees (GBDT) [27].
-We show that our models outperform baselines based on regressors or programming-language history.
-We consider the features defined in Tab.
-I.
-These features can be grouped into four categories.
-  1. 1.
-    Linguistic features: general properties characterizing a language pair like whether they support the same paradigms, the same style of memory management (is there a garbage collector or not), or the same type-system characteristics.
-  2. 2.
-    Syntactic features: properties of the syntax of a language pair measured by overlap in the counts of certain token types, as determined using the Pygments library on the corpus of code in the given languages.
-  3. 3.
-    Dataset-specific features: properties of the problems associated with code samples for the language pair in the online coding judge website from which the dataset came.
-  4. 4.
-    Model-specific features: whether the model saw the source language, target language, or both during pretraining.
-To predict the top source languages for a given task and target language $L_{t}$ in the set of target languages $T$, we train a ranker using the LightGBM [27] implementation of the LambdaRank algorithm [28].
-The model takes features in Tab.
-I as inputs and scores source languages $L_{s}$ in terms of their relevance to the target $L_{t}$.
+Details of the linguistic features are in Table LABEL:tab:linguistic-features.
+To predict the top source languages for a given task and target language $L_{t}$ in the set of target languages $T$, we train a ranker using the LightGBM (Ke et al., 2017) implementation of the LambdaRank algorithm (Burges, 2010).
+The model takes features in Table 9 as inputs and scores source languages $L_{s}$ in terms of their relevance to the target $L_{t}$.
 Our ranker uses a boosting ensemble of 100 decision trees with 16 leaves each.
@@
 We compare the performance of our ranker with a regression-based ranker and a history ranker.
+The regression-based ranker uses a LightGBM regressor with the same hyperparameter settings as our ranker and selects the top $K$ source languages with the highest predicted scores.
 The history ranker selects the top $K$ source languages with the shortest path to the given target language in a programming languages history graph [^2] http://rigaux.org/language-study/diagram.html (contracted to merge different versions of the same language).
-The regression-based ranker uses a LightGBM regressor with the same hyperparameter settings as our ranker and selects the top $K$ source languages with the highest predicted scores.
-Fig.
-6 shows the mean and standard deviation of NDCG@3 using LOO on $T$.
+Figure 7 shows the mean and standard deviation of NDCG@3 using LOO on $T$.
 While the error bars are overlapping, both regression and rankers outperform the history ranker on all tasks, indicating the importance of considering a variety of features for predicting top-performing sources rather than simplified heuristics.
@@
 Figure 7: NDCG@3 scores for different rankers and tasks corresponding to LOO evaluation over the set of target languages.
-Using a ranking model and considering a variety of features rather than simplified heuristics is crucial for predicting top performing sources.
-IV-C Feature analysis
-Figure 7: Normalized SHAP values aggregated by tasks for the features defined in Tab. I. The features are sorted by mean rank. Grayed-out features were unavailable for the given task.
-While Sec.
-IV-A explored how well transfer learning performs for different languages on different tasks, we would like to dig deeper into why they perform that way.
+The first four tasks are evaluated on ranking predictions on CodeT5 transfer experiments, while code generation is evaluated over ranking predictions on Llama 3.3 70B-Instruct.
+Rankers outperform regressors for predicting top-performing sources. Using historical relationships among programming languages alone is insufficient.
+4.3 Feature analysis
+Figure 8: Normalized SHAP values aggregated by tasks for the features defined in Table 9.
+The features are sorted by mean rank.
+Grayed-out features were unavailable for the given task.
+The first four tasks are evaluated on ranking predictions on CodeT5 transfer experiments, while code generation is evaluated over ranking predictions on Llama 3.3 70B-Instruct.
+While Section 4.1 explored how well transfer learning performs for different languages on different tasks, we would like to dig deeper into why they perform that way.
 That is, how do the features of language pairs affect transfer?
@@
 These questions can be answered by measuring feature importance.
-Unfortunately, most traditional techniques for measuring feature importance suffer from a common pitfall: when two features overlap, they arbitrarily assign more importance to one at the expense of the other.
-The state-of-the-art technique for measuring feature importance avoids this pitfall by considering all possible subsets $S\subseteq F$ of features [29].
+The state-of-the-art technique for measuring feature importance, SHAP values (Lundberg & Lee, 2017), considers all possible subsets $S\subseteq F$ of features.
 Then, the importance of a given feature is the difference it makes to model performance in combination with all possible sets of other features.
-The resulting SHAP values [29] show the relative impact of each feature on the model output.
 In the context of ranking different source languages for cross-lingual transfer, SHAP values measure how different features contribute to transfer.
-Fig.
-7 shows SHAP values computed based on the ranking models evaluated in Sec.
-IV-B.
+Figure 8 shows SHAP values computed based on the ranking models evaluated in Section 4.2.
 For presentation purposes, it rescales the values so the most important feature for each task has importance 1.
-Task-dependent feature importance
-While previous works emphasize the importance of specific features for cross-lingual transfer, the lack of comparison on different tasks limits their conclusions.
+Task-dependent feature importance.
+While previous works on transfer learning across programming languages emphasize the importance of specific features for cross-lingual transfer, the lack of comparison on different tasks limits their conclusions.
 The supported cross-task analysis demonstrates a task-dependent feature importance whereby different features contribute differently to cross-lingual transfer depending on the task.
@@
 A deeper understanding of the code semantics irrespective of the similar algorithms used is needed.
-The overlaps in names, keywords, and more generally tokens is key for understanding the semantics of code [11].
-Range of important features
-Different tasks seem to not only focus on different features, but also focus on a different number of features.
-While the feature importance for solution domain classification and error detection is concentrated on a small number of features, the feature importance is more spread out for clone detection and code repair.
-For example, the code repair task requires a transfer of knowledge from overlapping keywords, names, and more generally different tokens from a source language for fixing a code in a transfer language.
+The overlaps in names, keywords, and more generally tokens is key for understanding the semantics of code (Ahmed & Devanbu, 2022).
+Prior work on transfer learning in natural languages (Lin et al., 2019) showed feature importances to vary greatly across tasks. We confirm this for programming languages.
+Range of important features.
+Different tasks seem to not only focus on different features, but also on a different number of features.
+While the feature importance for solution domain classification and error detection is concentrated on a small number of features, it is more spread out for clone detection and code repair.
+For example, the code repair task requires a transfer of knowledge from overlapping keywords, names, and more generally different tokens from a source language for fixing a code in a target language.
 On the other hand, solution domain classification focuses on fewer features, related to the problems attributes.
-Comparison across categories
+Higher-level tasks seem to draw upon more features to best understand the code.
+Comparison across categories.
 Comparing the importance of different feature categories across tasks, we observe a relatively higher significance of syntactic features as demonstrated by the top two features Keywords and Names.
-Compared to transfer learning across natural languages [17], model-specific and linguistics features seem to carry less significance for transfer in programming languages.
+Compared to transfer learning across natural languages, model-specific and linguistics features seem to carry less significance for transfer in programming languages.
 While de Vries et al.
-[17] find the potential of cross-lingual transfer dropping significantly for sources and targets that are not seen by the model, our finding brings hope to not only transferring learning to low-resource target languages, but even finetuning the model on unseen languages that could serve as good sources, such as Kotlin.
-Within-category differences
+(2022) find the potential of cross-lingual transfer dropping significantly for sources and targets that are not seen by the model, our finding brings hope to not only transferring learning to low-resource target languages, but even finetuning the model on unseen languages that could serve as good sources, such as Kotlin.
+Overall, the most important features across the four tasks are keywords and names.
+Within-category differences.
 While different categories seem to have a higher impact on the model predictions, the impact of single features within a category varies across task.
-Our results are consistent with prior work that found Names and Keywords to be the most important features for cross-lingual transfer in certain tasks [11].
+Our results are consistent with prior work that found Names and Keywords to be the most important features for cross-lingual transfer in certain tasks (Ahmed & Devanbu, 2022).
 However, half of the tasks show keywords to be more important than names, while the other half has the opposite order of importance.
-1.
-The importance of language pair features for cross-lingual transfer performance depends on the task.
-2.
-Higher-level tasks seem to draw upon a broader range of features to best understand the code.
-3.
-Overall, the most important features across the four tasks studied in this paper are keywords and names.
-V Threats to Validity
-V-A External validity
-One threat is that to make the large number of experiments feasible and reduce their carbon footprint, we had to limit ourselves to a moderately-sized model.
-We mitigate this threat by picking CodeT5 [22], which is widely used in the research literature for AI-for-Code.
-We believe that even if we had used larger models, insights from one model would not perfectly generalize to another.
-Another threat is that to cover many languages, we restricted ourselves to four tasks.
-We mitigate this threat by picking tasks that cover a broad spectrum of difficulty and are diverse, as evidenced by their different performance and feature importances.
-Overall, our number of tasks exceeds prior work on cross-lingual transfer for programming languages.
-A further threat is that our datasets have some synthetic aspects, such as type-IV clones [25] for clone detection data or fault injection for code repair data.
-However, using fully natural data is such a high standard that hardly any academic paper satisfies it, nor is it typically the case in industry practice.
-Instead, we mitigate this threat by explicitly measuring the importance of dataset-specific features of language pairs.
-Another threat is that we focused on the setting of zero data in the target language.
+4.4 Few-shot prompting
+We examine cross-lingual transfer learning for programming languages by pairing low- and high-resource languages across classification and code-repair tasks.
+Whereas our initial study employed CodeT5-base, we now replicate a representative subset of experiments with the larger Llama 3.3 70B-Instruct model to test the generalizability of our conclusions to contemporary large-scale LLMs.
+For evaluation we adopt the few-shot prompting scheme described in Section 3; this technique has been shown to boost out-of-domain code-generation quality for previously unseen languages and to markedly reduce the compilation and parsing failures that arise when the model lacks explicit exposure to a target language (Athiwaratkun et al., 2023).
+We investigate whether rankers constructed from the features introduced earlier can accurately predict the best source language(s) for few-shot prompting.
+We compare different rankers explored in the previous section and test whether our ranker model can generalize to larger models.
+We train a ranker model on the same features described in Table 9 on Pass@1 scores.
+Figure 7 shows mean NDCG@3 scores.
+The learned ranker still surpasses other rankers and heuristic baselines, underscoring that considering multiple features is essential for accurately identifying the top source languages.
+SHAP analysis further reveals that expression of types and overlap in literals dominate the other features in predicting top performing sources.
+Languages whose type systems are richly expressed in code provide stronger semantic signals that transfer directly for code generation of other similarly typed languages solutions.
+Meanwhile a high lexical overlap in literals such as numeric and string constants offers immediate token-level anchors that reduce out-of-vocabulary risk during generation.
+Together, these two features jointly encode semantic structure and surface-level similarity, making them the most dependable indicators of best source languages for code generation.
+5 Limitations
+This section describes the limitations of our study, including the rationale for key design choices and the steps we took to address those limitations.
+5.1 External validity
+One limitation is that, to keep the large number of experiments feasible and reduce their carbon footprint, we constrained our fine-tuning experiments to a moderately sized model.
+We address this by using CodeT5 (Wang et al., 2021), a commonly adopted model in AI-for-Code research.
+Even if a larger model were used, insights from one model would not necessarily generalize to another.
+Accordingly, we included model-specific features in our study to account for potential model dependence.
+These features are described in Table 9 and analyzed as part of Section 4.3.
+Despite model-specific features, we do acknowledge that experimenting with different models, not just of varying size but also architecture, would have made our study more robust.
+Because the source and target languages are usually different and come in many combinations, we do not necessarily run the risk of overfitting.
+However, using models of varying strength could have exposed a broader range of more interesting patterns which would have made our study richer and more general.
+To mitigate this limitation, we conduct separate experiments with a larger model and few-shot prompting.
+Another limitation is that, to cover multiple languages, we limited ourselves to only five tasks.
+We address this by selecting tasks that span a broad range of difficulty and demonstrate diversity, as indicated by their differing performance and feature importance.
+Nonetheless, our study includes more tasks than any previous cross-lingual transfer study on programming languages.
+A further limitation is that our datasets include synthetic elements—for instance, type-IV clones (Roy & Cordy, 2007) in the clone detection data or fault injection in the code repair data.
+However, using exclusively natural data imposes a demanding constraint that few academic studies meet, and it is also uncommon in industrial practice.
+To address this, we explicitly measure the importance of dataset-specific features for each language pair.
+Another limitation is that we focused on the setting of zero data in the target language.
 This represents the lowest barrier to entry, as gathering even a small number of examples can be tricky.
-Therefore, we leave exploring the case where there is some target-language data to future work.
-V-B Internal validity
-We believe that our study avoids one common internal validity threat, which revolves around insufficient runs.
-Thanks to the sheer number of language pairs for each task, we get thousands of different experiments, whose results are consistent with each other.
-Moreover, each of these experiments involves fine-tuning with tens of thousands of samples, giving us confidence that the choice of samples is not a threat either.
-Since our study is the first of its kind at this scale, there is a threat that it cannot be directly compared to baselines.
-We mitigate this by observing that our results do confirm more narrow studies where applicable, for instance, a study of transfer for identifier-centric tasks [11].
-VI Conclusion
-We perform a systematic and extensive study of LLM transfer learning covering up to 41 programming languages, across 4 tasks including both classification and generation.
+Therefore, we defer exploring scenarios that include some target-language data to future work.
+5.2 Internal validity
+We believe that our study avoids one common internal validity limitation, which revolves around insufficient runs.
+Across all experiments, each source-target pair is evaluated only once.
+But thanks to the sheer number of language pairs for each task, we get thousands of different experiments, whose results are consistent with each other.
+Moreover, each of these experiments involves fine-tuning with tens of thousands of samples, giving us confidence that the choice of samples is not a limitation either.
+However, having results based on single runs does limit our ability to asses variability in experiments.
+Since our study is the first of its kind at this scale, it lacks directly comparable baselines.
+We mitigate this limitation by noting that, wherever comparisons are possible, our results corroborate those of narrower studies—for instance, prior work on transfer for identifier-centric tasks (Ahmed & Devanbu, 2022).
+6 Conclusion
+We perform a systematic and extensive study of LLM transfer learning covering up to 41 programming languages, across 5 tasks including both classification and generation.
+Programming languages covered include several low-resource but often still widely-used languages.
+Cross-lingual transfer learning performs much better than zero-shot with an LLM that has been pre-trained on code.
+One interesting finding is that even a language not seen during pre-training, like Kotlin, can be good fine-tuning source language across several target languages and tasks.
+On the other hand, certain languages that are often used extensively to pre-train LLMs, like C++ and Python, are worse source or target languages relative to others.
 To understand relative differences in cross-lingual performance between different programming languages, we define several linguistic, syntactic, dataset, and model-specific features of language pairs and then analyze the feature importance for a model that predicts transfer performance.
-Overall, we hope that this paper helps advance the community’s understanding of how learning transfers among programming languages, and that this improved understanding leads to better models to assist users of those languages, particularly low-resource ones.
-VII Data Availability
-The experiments are based on the publicly available CodeT5-base (220M parameters) model [22] and the open-sourced datasets CodeNet [19] and XCodeEval [20].
-Upon paper acceptance, we plan to open-source the code to create the derived datasets and the framework to train and evaluate the models.
-Also, we plan to make the 58 finetuned models available with the final version of the paper.
-We also plan to submit a replication package to the artifact committee.
+We show how different features are needed for ranking source languages, compared to ad-hoc heuristics such as overlap on names.
+To explain how a language that was unseen during pre-training, like Kotlin, can be a good source language, we show that seen language features are less significant compared to dataset, linguistic, and syntactic source language features.
+Similar to previous work of Ahmed & Devanbu (2022), we show that keywords and names are top features on average for CodeT5.
+On the other hand, unlike previous work, we cover more languages and more diverse tasks, and find that feature importance vary strongly across tasks.
+Replicating a representative subset of experiments with a larger model under few-shot prompting confirms that these feature hierarchies and the performance prediction gains they enable generalize to state-of-the-art LLMs.
+Overall, we believe that this paper sheds light on how learning transfers among programming languages, and that this ultimately leads to better models to assist users of those languages, particularly low-resource ones.
+7 Data Availability
+The experiments are based on the publicly available CodeT5-base (220M parameters) model (Wang et al., 2021) and the open-sourced datasets CodeNet (Puri et al., 2021) and XCodeEval (Khan et al., 2023).
+Our code is available at https://github.com/baltaci-r/XPL.